### v1

# Multi-critic Learning for Whole-body End-effector Twist Tracking

Abstract
Learning whole-body control for locomotion and arm motions in a single policy has challenges, as the two tasks have conflicting goals. For instance, efficient locomotion typically favors a horizontal base orientation, while end-effector tracking may require the base to tilt to extend reachability. Additionally, current Reinforcement Learning (RL) approaches using a pose-based task specification lack the ability to directly control the end-effector velocity, making smoothly executing trajectories very challenging. To address these limitations, we propose an RL-based framework that allows for dynamic, velocity-aware whole-body end-effector control. Our method introduces a multi-critic actor architecture that decouples the reward signals for locomotion and manipulation, simplifying reward tuning and allowing the policy to resolve task conflicts more effectively. Furthermore, we design a twist-based end-effector task formulation that can track both discrete poses and motion trajectories. We validate our approach through a set of simulation and hardware experiments using a quadruped robot equipped with a robotic arm. The resulting controller can simultaneously walk and move its end-effector and shows emergent whole-body behaviors, where the base assists the arm in extending the workspace, despite a lack of explicit formulations.
Figure 1: Smooth and precise whole-body loco-manipulation. We present a framework that enables quadrupedal robots to perform precise loco-manipulation through a multi-critic architecture and twist-based task space control. Our approach generates coordinated whole-body behaviors (Left) while maintaining accurate end-effector control during locomotion (Right).
Keywords: Loco-Manipulation, Multi-critic Reinforcement Learning, Whole-Body Control
1 Introduction
Robots need basic body control to perform manipulation tasks in various scenarios, such as industrial applications or service robotics. A robust low-level whole-body controller that can track body and end-effector poses and trajectories enables the seamless integration and execution of high-level tasks [1]. Multiple approaches to whole-body control exist, mainly divided between model-based control [2, 3, 4, 5], reinforcement learning (RL) based methods [6, 7, 1, 8, 9, 10, 11], and mixes of the two [12, 13, 14]. Model-based approaches have limitations and lack robustness due to the difficulty in accurately modeling interactions and the environment. RL has emerged as a staple framework for training locomotion controllers that can work in challenging environments [15, 16, 17, 18].
However, the field of whole-body control using RL has many challenges that are still being addressed. Firstly, locomotion and end-effector control have some conflicting goals. For example, locomotion tends towards a horizontal body orientation for an energy-efficient gait [15, 16, 17], but at the same time, tilting the base can increase the reachability of the arm. Similarly, velocity tracking of the base is far coarser than is needed to achieve precise end-effector control. Whole-body policies that can both locomote and move the arm tend to have low accuracy [7, 1]. Some solutions have attempted to either have separate base and arm controllers [12] or different policies for whole-body arm motion and locomotion [8]. However, these solutions all have limitations, either in requiring switching between controllers, not utilizing the entire potential workspace, or not being able to perform arm movements while in motion. A second challenge in whole-body control with RL is formulating the end-effector tracking task. Existing solutions use various ways to define goal pose tracking [13, 1, 10, 19, 8], or to perform end-effector force tracking [11]. However, trajectory following using a goal pose tracking controller may not always lead to smooth end-effector control since the controller attempts to rigidly track each control point. This is a fundamental issue with such task formulation due to its lack of velocity specification.
To address the first challenge of diverging goals, we propose using a multi-critic actor architecture [20, 21] that decouples the reward signals for locomotion, manipulation, and contact scheduling into individual reward groups. This simplifies the reward tuning, as separate critics allow the policy to resolve task conflicts more effectively. Furthermore, having separate critics seamlessly combines dense tracking rewards with sparse contact schedule rewards. We show that this approach is easier to use and also has desirable emergent behaviors.
As a second contribution, we design a 6D twist-based end-effector task formulation that enables direct control over the end-effector velocity while precisely tracking both discrete poses and continuous motion trajectories. Through comprehensive simulation and hardware experiments, we validate that the twist-based formulation leads to smooth end-effector tracking during stationary operation and during dynamic locomotion of the quadrupedal robot. In summary, our contributions are:
  • A multi-critic approach that trains a joint whole-body control policy that can simultaneously walk and accurately control the end-effector motion.
  • A twist-based formulation for whole-body end-effector tracking that allows for smoothly following any end-effector trajectory at varying velocities.
  • Experiments on simulation and hardware to demonstrate the effectiveness of our approach.
2 Related Work
Approaches towards whole-body control can be split into two major categories: learning-based or model-based control. Model predictive controllers (MPCs) can both locomote and perform high-accuracy end-effector motions [2, 3, 4], while avoiding self-collisions [5]. Although these methods can be further robustified by extending the model [22] or improving the optimization framework [23, 24], model-based approaches remain vulnerable to unmodeled disturbances. Especially for locomotion, reinforcement learning (RL) has emerged as the best technique to create robust and versatile controllers [15, 16, 17, 18], using large-scale parallel simulation [25, 26].
Similarly, for end-effector control, several RL-based controllers have been proposed [1, 6, 7, 8, 9, 10]. Some works tackle arm and base control as separate policies, where the arm is trained to reach a pose target and the base follows a separate velocity command. These methods either use RL for both [13], or a mix of model-based and RL control [14, 12]. Such split hierarchical approaches have limitations, as the arm cannot fully leverage the base policy to extend its reachability or improve accuracy by directly commanding the leg motors.
Conflicting goals also arise when training whole-body control policies. For example, the reward terms for locomotion tend to favor an overall horizontal base orientation while not enforcing a strict pitch or roll target, allowing for energy-efficient gaits to emerge [15, 16, 17, 18]. These controllers often do not even track small velocity changes. However, when training with end-effector accuracy in mind, a very stiff base is desired, which can also tilt to extend the reachability of the arm. Methods that attempt to balance this tuning problem, such as the ones presented by Fu et al. [7] or Ha et al. [1], usually suffer from accuracy or stability issues, as both approaches have end-effector tracking errors in the tens of centimeters. On the other hand, Portela et al. [8] split the problem into two separate whole-body controllers, one for walking and another for end-effector motions. Their approach can effectively tackle both the accuracy and robustness issues, but requires constant switching between controllers, and the arm can not be moved while the base is walking. In legged-wheeled systems, arm motions are more easily executed, as the presence of wheels reduces the necessity for base movements associated with walking gaits [9, 10]. Our proposed approach tackles this problem using a multi-critic approach [20, 21], enabling the locomotion and end-effector tracking problems to have separate sets of rewards, but one single policy. This emerging approach has been explored for style mixing [20] or locomotion [27, 28]. We showcase further uses and how multi-critic RL training simplifies the tuning problem and emergently leads to desired behaviors.
Another critical aspect of using RL for end-effector tracking is the target and reward formulation for arm movement. Early works focused solely on 3D position tracking [7, 6], which later evolved into full 6D pose tracking, using a variety of representations, from Euler angles [13], to rotation matrices [1, 10, 19], and keypoints [8]. However, for pose-based approaches, smoothly tracking a trajectory can be difficult and requires a high-level policy to monitor the execution [1, 9]. An alternative to pose tracking is force tracking, as demonstrated by Portela et al. [11], which allows the possibility of building an impedance controller on top. In contrast, our approach circumvents these issues by directly performing twist tracking, resulting in simpler reward tuning for the entire problem. Furthermore, the twist tracking formulation provides explicit control over end-effector velocity, enabling the robot to follow complex trajectories.
3 Method
Figure 2: Architecture for the teacher training pipeline: Given a randomly sampled start and goal pose, the desired end-effector and base twist, along with the desired foot height, are provided as commands to the policy by the command generator. Rewards are categorically computed and consumed by separate critics, leading to individual value functions. The advantage is estimated per critic, normalized, and summed to compute the total advantage. Policy optimization is performed using Proximal Policy Optimization (PPO) to optimize the teacher policy.
We aim to train a controller that takes locomotion and end-effector task commands and outputs joint position residuals, enabling coordinated whole-body motion. More specifically, we want the ability to smoothly track 6D motion trajectories with explicit control over the end-effector velocity. Lastly, we want to ensure that torso motions do not affect the manipulation task and that the robot can perform the manipulation task while stationary and during walking.
3.1 Task definition
We formulate the task as a Partially Observable Markov Decision Process (POMDP) and train the policy using a teacher-student architecture (Fig. 2) that can follow a target command $\bm{c}_{t}$ defined as
$\bm{c}_{t}=[_{B}\bm{v}_{b},~{}_{B}\bm{\omega}_{b},~{}_{B}\bm{v}_{EE},~{}_{B}\bm{\omega}_{EE},~{}_{B}\bm{\hat{p}}_{G},~{}\bm{\hat{R}}_{BG},~{}\bm{h}_{f}],$ (1)
where ${}_{B}\bm{v}_{b}$ and ${}_{B}\bm{\omega}_{b}$ are the base linear and angular velocity targets expressed in robot’s base frame, ${}_{B}\bm{v}_{EE}$ and ${}_{B}\bm{\omega}_{EE}$ are the end-effector linear and angular velocity targets expressed in robot’s base frame, ${}_{B}\bm{\hat{p}}_{G}$ and $\bm{\hat{R}}_{BG}$ define the final pose in the command trajectory expressed in robot’s base frame, and $\bm{h}_{f}$ is the desired swing heights for each foot.
At each control time step $t$, the teacher policy $\pi_{\theta}^{T}$ generates an action $\bm{a}_{t}$ taking as input $\bm{s}_{t}~{}=~{}[\bm{h}_{t},~{}\bm{o}_{t},~{}\bm{p}_{t},~{}\bm{a}_{t-1},~{}\bm{c}_{t}]\in\mathbb{R}^{187}$ consisting of the 4 step history of previous joint positions $\bm{h}_{t}\in\mathbb{R}^{72}$, a set of observations concerning the robot state $\bm{o}_{t}\in\mathbb{R}^{45}$, privilege information obtained from the simulator $\bm{p}_{t}\in\mathbb{R}^{32}$, the action at previous time step $\bm{a}_{t-1}\in\mathbb{R}^{18}$, and the command $\bm{c}_{t}\in\mathbb{R}^{20}$. For more details, please refer to the Appendix. We normalize the observations using a running buffer to ensure that the policy can learn a robust representation of the state space.
Given the observations, the policy $\pi_{\theta}^{T}$ generates an action $\bm{a}_{t}=[\bm{a}_{t}^{arm},~{}\bm{a}_{t}^{robot}]$ where $\bm{a}_{t}^{arm}\in\mathbb{R}^{6}$ and $\bm{a}_{t}^{robot}\in\mathbb{R}^{12}$ are the relative offsets to be applied. Unlike the standard approach[8] of applying actions to predefined default joint positions, we apply these offsets directly to the current joint positions. Assuming a Gaussian action distribution, the actions are normalized with a fixed standard deviation $\sigma_{a}$ and clamped to be within $[-1,1]$. The joint position targets for the next time step $\bm{q}_{t+1}$ are computed as,
$\bm{q}_{t+1}=\bm{q}_{t}+\text{clamp}((\bm{a}_{t}/\sigma_{a}),~{}-1,~{}~{}1)$ (2)
3.2 Command Formulation
End-effector trajectory sampling: During training, we generate manipulation trajectories by initializing the start pose $\bm{T}_{s}$ at the current end-effector location, while the goal pose $\bm{T}_{g}$ is randomly sampled using a two-stage approach. First, we sample positions within a unit sphere centered at the robot’s shoulder. Any samples falling within a predefined cuboid representing the robot’s torso and hip region are rejected to ensure physically feasible goals, which enables smoother training [8]. For orientation determination, we construct a reference orientation where the x-axis aligns with gravity and the z-axis points along the vector from the torso center to the sampled position. This reference orientation is then perturbed with random rotations about each axis, bounded within ±$\textpm\frac{\pi}{4}$ radians. This approach ensures that a wide range of the sampled poses are within the robot’s feasible workspace. Given the start and goal pose, we implement a temporal trajectory formulation that prioritizes continuous tracking rather than terminal goal achievement. We assign a random duration to each trajectory and generate intermediate waypoints to be tracked using linear interpolation in position and spherical linear interpolation in orientation. The intermediate goals are defined as
$\bm{r}_{i}$ $=\bm{r}_{s}+\alpha\cdot(\bm{r}_{g}-\bm{r}_{s})$ $\bm{r}$ $\in\mathbb{R}^{3}$ (3)
$\bm{\theta}_{i}$ $=\bm{\theta}_{s}\boxplus\alpha\cdot(\bm{\theta}_{g}\boxminus\bm{\theta}_{s})$ $\bm{\theta}$ $\in SO(3)$
[…]
End-effector command: While most pose tracking methods in the literature [13, 1, 10, 19, 8] define the task to track a final goal pose within a given time, our task formulation is a continuous trajectory tracking problem with explicit velocity control. Hierarchical approaches, which send intermediate goal poses to a pose-based tracking controller to achieve trajectories, lack critical information about the required end-effector velocity. This deficiency leads to jerky end-effector motions as the policy attempts to rigidly maintain each intermediate pose without smoothly transitioning between them. To address these limitations, we introduce an end-effector twist-based command formulation that explicitly incorporates velocity information. More specifically, we define the command as:
$\bm{v}_{EE}=\frac{1}{\Delta t}(\bm{r}_{i}-\bm{r}_{EE})\ \ \text{and}\ \ \bm{\omega}_{EE}=\frac{1}{\Delta t}(\bm{\theta}_{i}\boxminus\bm{\theta}_{EE})$ (4)
where, $\bm{r}_{EE}$ and $\bm{\theta}_{EE}$ are the current end-effector position and orientation respectively, and $\Delta t$ is the control time step. Finally, the end-effector command to the policy is defined as the stacked vector of $\bm{v}_{EE}$, $\bm{\omega}_{EE}$, and the final goal pose $\bm{T}_{g}$ along the entire trajectory.
We observe that training exclusively with global feed-forward trajectories as defined above can lead to a sparse distribution of positive tracking examples, leading to extended training times before the policy grasps the task. To mitigate this, we define a subset of the trajectories locally. A local trajectory is created by re-initializing the start pose of the trajectory at every control time step. This approach ensures that the policy frequently encounters states near the desired trajectory, providing a denser representation of the reward landscape and accelerating the learning process. Similarly to [8], we represent the end-effector trajectory in a robot-centric task frame that is agnostic to the torso height, pitch, and roll orientation. This choice decouples the task from the torso motion, allowing the policy to learn a coordinated whole-body motion to satisfy a precise end-effector motion.
Foot trajectory: Foot trajectory generation computes the desired foot heights above the ground during locomotion. To achieve this, a gait phase $\phi$ and foot phase offsets $\bm{\theta}=[\theta_{LF},~{}\theta_{RF},~{}\theta_{LH},~{}\theta_{RH}]$ are used to generate a sinusoidal swing foot trajectory pattern [14], where the subscripts denote the four legs: Left Front (LF), Right Front (RF), Left Hind (LH), and Right Hind (RH) . The height of the feet is computed as:
$\begin{split}\bm{h}_{f}&=h_{max}[\text{sin}(2\pi\theta_{LF}),~{}\text{sin}(2\pi\theta_{RF}),~{}\text{sin}(2\pi\theta_{LH}),~{}\text{sin}(2\pi\theta_{RH})]\\ \bm{\theta}&=[\phi+\theta_{LF},~{}\phi+\theta_{RF},~{}\phi+\theta_{LH},~{}\phi+\theta_{RH}]\end{split}$ (5)
During an episode, the gait phase is updated according to a fixed gait frequency, and the foot phase offsets are fixed to correspond to a static gait pattern.
Base command: The base command consists of the linear velocity of the torso ${}_{B}\bm{v}_{b}=[_{B}\bm{v}_{b_{x}},_{B}\bm{v}_{b_{y}}]$ and the angular velocity ${}_{B}\bm{\omega}_{b}$ along the Z-axis wrt. the base frame. During training, this velocity command is periodically resampled during the episode from a uniform distribution within the range $[-0.25,0.25]$.
3.3 Multi-critic Actor Learning
Multi-critic actor learning has previously been shown to be effective in multi-task RL problems [20] and to effectively manage mixtures of dense and sparse rewards [27]. Fu et al. [7] introduced two critic networks and used curriculum learning to gradually mix two advantage functions to combine the learning of arm and leg actions in a unified policy. In this work, we categorize rewards into three distinct groups: locomotion, manipulation, and foot contact schedule. For each group, we define a dedicated critic $\phi_{i}$ that estimates the corresponding value function $V_{\phi_{i}}(.)$. The overall advantage $\hat{A}$ is calculated as the sum of the normalized advantage estimates $\hat{A}_{i}$ from each critic. Following the PPO framework, we compute the policy surrogate loss as
$L(\theta)=\hat{\mathbb{E}}[min(r_{t}(\theta)\hat{A},~{}clip(r_{t}(\theta),~{}1-\epsilon,~{}1+\epsilon)\hat{A})]$
where $r_{t}(\theta)$ is the ratio of new to the old policy probabilities, $\epsilon$ is the PPO clipping parameter, and $\hat{\mathbb{E}}$ is the empirical average over the batch. The three reward groups are structured as follows:
  • The locomotion group encompasses base velocity tracking rewards and regularization rewards for locomotion actions.
  • The manipulation group includes end-effector position and orientation tracking rewards, along with associated regularization terms.
  • The contact schedule group comprises rewards for tracking foot contact states and corresponding regularization terms.
Notably, we introduced the contact schedule as a separate group due to the low-frequency nature of gait patterns, which results in a relatively sparse reward distribution compared to the other groups.
3.4 Teacher-Student Distillation
Our teacher policy is a 3-layer MLP with hidden layers of size $[512,256,128]$ with ELU activations. The teacher has access to privileged information inaccessible in the real world. To deploy the policy on hardware, a student policy with the same network structure as the teacher is distilled without access to privileged information using supervised learning [16].
4 Results
We deploy our controller on the ANYmal D robot [29] from ANYbotics AG mounted with a Dynaarm manipulator from Duatic AG. The robot has a total of 18 actuators, of which 12 actuate the legs. While ANYmal’s motion controller is executed at 400 Hz, our control policy is executed with a decimation rate of 8:1, resulting in 50 Hz on the onboard computer [8, 17, 30].
4.1 Hardware Experiments
Figure 3: End-Effector Tracking Performance for Different Trajectories: Desired and measured position and velocity for tracking linear, circular, and semicircular trajectories (left to right).
Tracking performance of different trajectory types: To assess the tracking performance of the policy, we evaluate its ability to track the following trajectory types:
  • Straight line: The policy is commanded to move 30 cm along each axis.
  • Circle: The policy is commanded to move in a circle on the YZ plane with a radius of 20 cm at a fixed distance in front of the robot.
  • Semicircle around robot: To evaluate trajectories over an extended workspace, the policy is commanded a trajectory of radius 60 cm radius around the robot.
During these experiments, the end effector orientation is actively controlled to point outwards along the vector from the torso center to the end effector while aligning the X-axis of the end-effector frame with gravity vector. In Fig. 3, we show the tracking performance of the policy for all three trajectory types.
Tracking trajectories with varying velocities:
Motion $\hat{\bm{v}}_{EE}$ $\delta\bm{r}$ $\delta\dot{\bm{r}}$
Linear 0.05 0.0166 0.0519
0.1 0.0175 0.0623
Circular 0.05 0.0220 0.0609
0.1 0.0274 0.0898
0.2 0.0348 0.1781
Semicircular 0.1 0.0480 0.2339
0.2 0.0732 0.3590
Table 1: Tracking performance in end-effector position error (L2 norm) $\delta\bm{r}$ (m) and velocity error $\delta\dot{\bm{r}}$ (m/s) for robot linear motions XYZ executed at specified distances and speeds $\hat{\bm{v}}_{EE}$ (m/s).
To demonstrate the capability of the controller to execute trajectories at varying velocities, we execute the three trajectory types described above at different commanded velocities. The RMSE end-effector tracking error in position and velocity while tracking trajectories at different speeds is listed in Table 1. As also shown in Fig. 3, the controller achieves good velocity tracking across different velocity ranges. Although velocity tracking accuracy is somewhat lower at slower speeds, particularly for linear trajectories, the policy successfully maintains low position errors throughout the execution. It must be noted that circular and semicircular trajectories being evaluated are not part of the training distribution, since the trajectories in training are always computed between start and goal pose through linear interpolation.
Chicken head control performance while walking:
Figure 4: End-Effector Tracking performance in chicken-head mode: Tracking error in end effector position (red, left axis) and measured base velocity (blue, right axis) of the robot plotted over time. The position error has a mean of 0.0156 m and maximum of 0.0287 m.
To demonstrate the whole-body task coordination capabilities of the controller, we perform chicken head control with the end-effector while walking. The goal of this experiment is to demonstrate the whole-body loco-manipulation capability of the policy by holding a static end-effector pose in the world frame while moving the torso. During this experiment, the robot is commanded a base velocity within the range of [-0.2, 0.2] m/s while the end-effector is commanded twist corrections to hold the end-effector position in the robot’s odometry frame. As shown in Fig. 4, the policy achieves very low end-effector position tracking errors while walking, demonstrating effective task space control during base motion.
4.2 Simulation Experiments
Effect of multi-critic architecture on gait tracking: Multi-critic architectures have previously demonstrated effectiveness in multi-task RL problems. While the agent is trained only with a static walking gait pattern, we observe that the policy generalizes to perform a trotting gait at runtime by simply modifying the commanded foot swing trajectories. The measured foot trajectories for both static walk and trotting gaits are shown in Fig. 5.
Figure 5: Contact Schedule Adaptation to Unseen Gait Patterns: Measured height of Left Fore (LF), Right Fore(RF), Left Hind (LH) and Right Hind (RH) feet for (a) static walking measured on hardware, and (b) trot gait measured in simulation.
This result demonstrates that separating the contact schedule into a separate critic enables the policy to learn a more general representation of gait patterns, allowing it to adapt foot timings based on the input command. We note that such generalization has not been previously reported in the literature, highlighting the efficacy of our multi-critic architecture.
Comparison with existing methods: To evaluate our proposed approach against existing methods, we conduct comparative experiments using a semicircular end-effector trajectory, both while the robot is stationary and during locomotion. We integrate our arm control with three different locomotion policies: Ma [12], Taka 3-DoF [17], and Taka 6-DoF [30]. To achieve this, we apply the leg actions generated by the respective locomotion policy while executing the arm actions produced by our proposed approach. We compute the tracking end effector tracking errors using the measurements obtained from the state estimator. As shown in Fig. 6, while all combinations achieve satisfactory end-effector tracking, our proposed whole-body locomotion policy demonstrates superior performance with the lowest tracking errors.
State Controller $\delta\bm{r}$ [m] $\delta\bm{\theta}$ []
Standing Proposed 0.0176 1.815
Taka 3-DoF [17] 0.0580 4.196
Taka 6-DoF [30] 0.0439 3.608
Ma [12] 0.0386 3.515
Portela [8] 0.0391 1.756
Walking Proposed 0.0358 2.872
Taka 3-DoF [17] 0.4180 4.711
Taka 6-DoF [30] 0.0436 4.974
Ma [12] 0.0483 6.432
Portela [8] - -
Figure 6: Average end-effector tracking errors: $\delta\bm{r}$ (mean Euclidian norm) for position and $\delta\bm{\theta}$ (mean angular error) for orientation, across controllers while standing and walking.
For comparison with existing pose-based controllers, we evaluate against the whole-body policy from Portela et al. [8]. Since their policy is not designed for locomotion, we present results only for stationary experiments. Our approach achieves significantly better task space trajectory tracking in position while maintaining comparable orientation accuracy. Notably, stationary policies typically have a substantial advantage during training due to enhanced stability from continuous leg contact. This makes our results particularly significant, as our proposed approach delivers superior performance while being able to perform loco-manipulation.
5 Conclusion
In this work, we presented a whole-body loco-manipulation framework for quadrupedal robots that enables continuous and smooth trajectory tracking while stationary and during locomotion. Through extensive simulations and hardware experiments, we demonstrated the effectiveness of our twist-based command formulation in achieving precise end-effector velocity control. Our novel multi-critic architecture successfully combines locomotion, manipulation, and contact scheduling without requiring complex training curricula or reward tuning. Comparative evaluations against existing non-whole-body and pose-based controllers validate the superior performance of our approach.
6 Limitations
The current implementation has two main limitations. First, locomotion performance is restricted to moderately rough terrains, as robust rough-terrain controllers were not the focus of this work. In future work, we plan to address this by including reward formulations and perception mechanisms from existing rough-terrain locomotion frameworks. Our multi-critic formulation should greatly simplify the tuning involved in incorporating these new terms. However, having to train the whole-body teacher to incorporate both robust arm maneuvers and perceptive locomotion will significantly increase training times and make later tuning of individual behaviors more difficult. The second limitation is that during whole-body maneuvers, the robot occasionally reaches the limits of the shoulder joint, triggering safety mechanisms. In the future, this can be addressed by incorporating joint limits during policy training, either as penalties, terminations, or, for example, through the use of constrained PPO.